Q.2 Player 1 and player 2 bargain over sharing 300 dollars. The asymmetric Nash product is: 2- (x1 - 20)1/3(x2 10)2/3. Find the bargaining solution.
Answer #1

Answer:

As given

X1+X2=300

X1>0

X2>0

Nash product =(X1-20)^(1/3)*(X2-10)^(2/3)

Nash product = (X1-20)^(1/3)*(300-X1-10)^(2/3)

Nash Product=(X1-20)^(1/3)*(290-X1)^(2/3)

For Bargaining solution d(Nash product)/dX1=0

(1/3)*{(290-X1)/(X1-20)}^(2/3)-(2/3)*{(X1-20)/(290-X1)}^(1/3)=0

290-X1=2*(X1-20)

X1=$110

X2=300-110=$190